PEGCOR has a​ P/E ratio of 15.0. Earnings per share are $2.00​, and the expected EPS 5 years from today are $3.22.

Calculate the PEG ratio.​
Answer #1

PEG ratio=PE/growth rate=PE/100*1/((EPS 5 years from today/EPS now)^(1/5)-1)=15/100*1/((3.22/2)^(1/5)-1)=1.501045872